OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form. –x + 4y = –1 x – 4y = –1 x + y = 3 x – y = –3

OpenStudy (blurbendy):

y -4 = -1(x + 1) y = -x -1 + 4 y = -x + 3